Which of the following best describes average fixed cost.

a.

Variable costs times number of units produced

b.

ATC minus AVC

c.

Total costs divided by number of units produced

d.

All describe averaged fixed costs
Answer #1

Total cost is the sum of fixed cost and variable cost which means

TC = FC + VC.

Divide the equation by Q on both sides

This says that

TC/Q = FC/Q. + VC /Q

when divided by Q total becomes average which means

ATC = AFC + AVC

ATC - AVC = AFC

Therefore (b) is the answer